OpenStudy (anonymous):

Simplify. Express the quotient as a rational exponent. x^7/8/x^2/5

OpenStudy (anonymous):

5x^5/8

OpenStudy (anonymous):

not an answer choice.

OpenStudy (anonymous):

x^(19/40)

OpenStudy (anonymous):

\[\frac{\frac{x^7}{8}}{\frac{x^2}{5}}\] before i thought u asked for this expression

OpenStudy (anonymous):

\[\Large \begin{array}{l} \frac{{{x^{7/8}}}}{{{x^{2/5}}}}\\ = {x^{(7/8) - (2/5)}}\\ \frac{7}{8} - \frac{2}{5}\\ = \frac{{35}}{{40}} - \frac{{16}}{{40}}\\ = \frac{{19}}{{40}}\\\\ \text{Hence:}\\ {x^{19/40}} \end{array}\]